# Shared Global and Local Geometry of Language Model Embeddings

### v2

@@ Abstract
Researchers have recently suggested that models share common representations. In our work, we find that token embeddings of language models exhibit common geometric structure. First, we find “global” similarities: token embeddings often share similar relative orientations. Next, we characterize local geometry in two ways: (1) by using Locally Linear Embeddings, and (2) by defining a simple measure for the intrinsic dimension of each token embedding. Our intrinsic dimension demonstrates that token embeddings lie on a lower dimensional manifold. We qualitatively show that tokens with lower intrinsic dimensions often have semantically coherent clusters, while those with higher intrinsic dimensions do not. Both characterizations allow us to find similarities in the local geometry of token embeddings. Perhaps most surprisingly, we find that alignment in token embeddings persists through the hidden states of language models, allowing us to develop an application for interpretability. Namely, we introduce Emb2Emb, a simple method to transfer steering vectors from one language model to another, despite the two models having different dimensions.
@@ 1 Introduction
To this point, we study the similarities in geometric properties of the embedding space of language models. Although alignment across word embeddings have been studied before, prior work has mostly been across embedding techniques (e.g., GloVe (Pennington et al., 2014) vs. word2vec (Mikolov et al., 2013b)(Dev et al., 2019), across modalities (Huh et al., 2024; Merullo et al., ), or in cross-lingual settings (Mikolov et al., 2013a; Alvarez-Melis & Jaakkola, 2018; Artetxe et al., 2017; Conneau et al., 2020). In our work we characterize the geometry of token embeddings in contemporary large language models and study the similarities across language models.
@@
Perhaps most surprising is that the alignment in token embeddings seem to persist in the hidden layers of the language models. Given these insights, we introduce Emb2Emb, a simple tool for model interpretability: steering vectors that can control one model can be transferred and reused for another model, despite the models having different dimensions.
@@ 2 Related Work
Shared Representations.
Researchers suggest that neural networks share common representations. Bansal et al. (2021) study representation similarity by “stitching” (Lenc & Vedaldi, 2015) the early layers of one model with the later layers of another and showing minimal change in performance. More recently, Huh et al. (2024) pose the Platonic Representation Hypothesis: large models across different modalities are converging towards the same representations, given the vast amounts of training data that are used by these models.
Our work adds evidence of similar representations across models. Namely, we show that embeddings of language models share similar geometric structures.
Input Embeddings.
Input embeddings are often the “backbone” of neural networks. Most recent models, including Transformers, use residual connections (He et al., 2016), meaning that the input embeddings are the start of the “residual stream” (Elhage et al., 2021), to which subsequent layers iteratively construct features (Jastrzebski et al., 2017).
Such embeddings often encode vast amounts of information. Word2vec (Mikolov et al., 2013b) famously demonstrated that relational information of tokens may be linearly represented. More recently, Park et al. (2024a) demonstrate how hierarchical information is encoded in the token embeddings of contemporary language models.
@@
We study the embeddings of contemporary large language models, and find that aligning embeddings allows for applications such as transferring steering vectors across models.
@@
Researchers have found that often, language models use linear representations for various concepts (Nanda et al., 2023; Li et al., 2024; Park et al., 2024b; Rimsky et al., 2023). Conveniently, linear representations suggest that such representations can be easily manipulated with simple vector arithmetics to control the model’s behavior. Researchers have referred to such vectors that can play a causal role as “steering vectors” (Turner et al., 2023). In our work we demonstrate that given the similar geometric properties of models’ embeddings, steering vectors can be transferred from one language model to another.
@@ 3 Preliminary
We briefly review the Transformer architecture to set notation.
The forward pass starts with assigning each input token an embedding from a learned embedding matrix $\mathcal{E}\in{\rm I\!R}^{V\times d}$, where $V$ is the vocabulary size and $d$ is the embedding dimension. The token embeddings form the initial hidden states, denoted as $\mathbf{h}^{0}$.
@@
In our work we consider language models from the same “family”, which each have different sizes (i.e., embedding dimension, number of layers, etc.), but share the same tokenizer. [^1] Unfortunately, we lack information about the training data used to train each model, but we assume a large overlap in the training data used within each family. We study three families: GPT2 (small, medium, large, xl), Llama3 (1B, 3B, 8B, 11B-Vision, 70B), and Gemma2 (2B, 9B, 27B) (Radford et al., 2019; Dubey et al., 2024; Team et al., 2024).
@@
Figure 1 shows our results, with most language models demonstrating high Pearson correlations. These results indicate that most language models within the same family share similar relative orientations of token embeddings. Note that this is despite the fact that the language models have different embedding dimensions. Further note that for Llama3, the base models and their instruction-tuned counterparts always have a correlation scores near 1, indicating that the global geometry of token embeddings does not change much. Lastly, in the case of untied token embeddings (Llama3 8B, 11B-V, 70B), we interestingly only observe high correlation scores for the unembedding space, but not for the embedding space. We leave further exploration of the impact of tied versus untied embeddings for future work.
@@ 5.1 Locally Linear Embeddings of Tokens
Figure 2: Language models share similar local structure. We compare LLE weights of token embeddings across language models, where LLE weights capture the local structure of each embedding. Most token embeddings have very similar local structure across language models, as indicated by the density accumulated around high cosine similarity scores.
@@
Note that some tokens have a cosine similarity of 0. These may be undertrained tokens (Land & Bartolo, 2024) – tokens present in the tokenizer, but not in the training data (e.g., “_SolidGoldMagikarp”). Namely, their embeddings are likely random weights assigned during initialization. Note that in high dimensions, random vectors are likely to be orthogonal.
@@ 5.2 Measuring Intrinsic Dimension
We also characterize the local geometry of embeddings by defining a simple measure for intrinsic dimension (ID). To measure the ID of a token, we use $k$ (= 1,200) of its nearest neighbors. We then run PCA on the $k$ points, and refer to the number of principal components needed to explain some threshold amount (95%) of the variance as the token’s intrinsic dimension. Varying hyperparameter values for $k$ ([500, 1,200, 2,000]) and variance threshold ([0.8, 0.9, 0.95]) yield similar trends: while the absolute ID values differ, tokens with lower ID yield semantically coherent clusters.
@@ 5.3 Low Intrinsic Dimensions Indicate Semantically Coherent Clusters
Qualitatively, tokens with lower IDs exhibit semantically similar tokens, while tokens with higher ID show less coherence. We also show the ID of a random vector as a baseline. While further analysis is needed, roughly, tokens with semantic clusters seem to be below the random baseline while those with less coherence are above the baseline ID. What determines the ID of tokens? We view this as an exciting question and leave it for future work.
@@ 6 Emb2Emb: Transferring Steering Vectors
Our setup for Llama3 is the same as that of Rimsky et al. (2023). We take human-written evaluation datasets from Perez et al. (2022) and Rimsky et al. (2023), which contain questions with two answer choices. One choice answers the question in a way that demonstrates the target behavior, while the other does not. Examples can be found in Appendix C.
@@ Results.
Figure 4 demonstrates steering Llama3-8B with steering vectors transferred from Llama3-1B and 3B (See Appendix D for more examples of transferred steering). In each subplot, the dotted curve indicates a point of reference in which we steer the target model using a steering vector from the same model (i.e., $\mathbf{v}_{S}$ == $\mathbf{v}_{T}$), while solid lines indicate a steering vector transferred from a different model. The legends indicate the source model and layer from which a steering vector is taken from, as well as the layer in the target model that is intervened on. The x-axis indicates how much each steering vector has been scaled ($\alpha$ of Equation 5), while the y-axis indicates the likelihood of the model choosing an option that exhibits a behavior of interest. In most cases, the transferred steering vectors exhibit similar trends as the original steering vector.
@@ 7 Discussion
Our findings add evidence that similar representations can be found across models. Namely, we show that the token embedding space of language models share similar global and local geometric structures.
First, we find “global” similarity: token embeddings have similar relative orientations. Second, we find that token embeddings also share local geometric structures. We demonstrate this by looking at LLE weights of tokens, as well as by defining a simple measure for intrinsic dimension. Our ID measure reveals an interesting pattern, where tokens with lower intrinsic dimensions form semantically coherent clusters.
Perhaps most surprisingly, alignment in the model’s embedding space seem to persist through the hidden states. This leads to Emb2Emb, a simple application for model interpretability: steering vectors can be transferred across language models.
@@ Appendix A Closed Form Solution for LLE
Appendix B Additional Results for LLE Similarity
Here we show results from comparing Locally Linear Embedding weights across language models. Figure 6, 7, and 8 demonstrate results for GPT2, Llama3 (embeedings), and Gemma2.
@@ Appendix B Additional Results for LLE Similarity
Appendix C Example of Steering Data
@@
Appendix D Additional Examples of Transferred Steering for Llama3

### v3

@@ Abstract
Researchers have recently suggested that models share common representations. In our work, we find numerous geometric similarities across the token embeddings of large language models. First, we find “global” similarities: token embeddings often share similar relative orientations. Next, we characterize local geometry in two ways: (1) by using Locally Linear Embeddings, and (2) by defining a simple measure for the intrinsic dimension of each embedding. Both characterizations allow us to find local similarities across token embeddings. Additionally, our intrinsic dimension demonstrates that embeddings lie on a lower dimensional manifold, and that tokens with lower intrinsic dimensions often have semantically coherent clusters, while those with higher intrinsic dimensions do not. Based on our findings, we introduce Emb2Emb, a simple application to linearly transform steering vectors from one language model to another, despite the two models having different dimensions.
@@ 1 Introduction
To this point, we study the similarities in geometric properties of the embedding space of language models. Although alignment across word embeddings have been studied before, prior work has mostly been across embedding techniques (Dev et al., 2019) (e.g., GloVe (Pennington et al., 2014) vs. word2vec (Mikolov et al., 2013b)), across modalities (Huh et al., 2024; Merullo et al., ), or in cross-lingual settings (Mikolov et al., 2013a; Alvarez-Melis & Jaakkola, 2018; Artetxe et al., 2017; Conneau et al., 2020). In our work we characterize the geometry of token embeddings in contemporary large language models and study the similarities across language models.
@@
Perhaps most surprising is that the alignment in token embeddings seem to persist in the hidden layers of the language models. With these insights, we introduce Emb2Emb, a simple tool for model interpretability: steering vectors that can control one model can be linearly transformed and reused for another model, despite the models having different dimensions.
@@ 2 Related Work
Shared Representations & Geometry.
Researchers have studied the geometry of language model representations in the past (Valeriani et al., 2023; Doimo et al., 2024). Interestingly, researchers suggest that neural networks often share common representations. Bansal et al. (2021) study representation similarity by “stitching” (Lenc & Vedaldi, 2015) layers across models, with minimal change in performance. More recently, Huh et al. (2024) pose the Platonic Representation Hypothesis: large models across different modalities are converging towards the same representations, given the vast amounts of training data that are used by these models. In our work we demonstrate numerous similarities in the geometry of language model embeddings.
Geometry of Embeddings.
Embeddings are often the “backbone” of neural networks. Most recent models, including Transformers, use residual connections (He et al., 2016), meaning that the input embeddings are the start of the “residual stream” (Elhage et al., 2021), to which subsequent layers iteratively construct features (Jastrzebski et al., 2017).
Such embeddings often encode vast amounts of information. Word2vec (Mikolov et al., 2013b) famously demonstrate that relational information of tokens may be linearly represented. More recently, Park et al. (2024a) demonstrate how hierarchical information is encoded in the token embeddings of contemporary language models.
Researchers have studied the geometry of embeddings before. Burdick et al. (2021); Wendlandt et al. (2018) study the instability of word embeddings such as word2vec or GloVe. Unlike their work, we find that the embeddings of contemporary language models share numerous geometric similarities.
Papyan et al. (2020) discover neural collapse, in which the penultimate layer and class representations of a network converge to a simplex equiangular tight frame. Zhao et al. (2024) provides a theoretical understanding of the implicit geometry that results from next-token prediction. Our work is closely related to both lines of work, and may be an empirical instantiation of their findings.
@@
Researchers have found that often, language models use linear representations for various concepts (Nanda et al., 2023; Li et al., 2024; Park et al., 2024b; Rimsky et al., 2023). Conveniently, linear representations allows one to easily manipulate the representations with simple vector arithmetics to control the model’s behavior. Researchers have referred to such vectors as “steering vectors” (Turner et al., 2023). We demonstrate that given the similar geometry, steering vectors can be linearly transformed from one language model to another, also demonstrated by concurrent work from Oozeer et al. (2025).
@@ 3 Preliminary
We briefly review the Transformer architecture to set notation. The forward pass starts by assigning each input token an embedding from a learned embedding matrix $\mathcal{E}\in{\rm I\!R}^{V\times d}$, where $V$ is the vocabulary size and $d$ is the embedding dimension. The token embeddings form the initial hidden states, denoted as $\mathbf{h}^{0}$.
@@
In our work we consider language models from the same “family”, which each have different sizes (i.e., embedding dimension, number of layers, etc.), but share the same tokenizer.We study three families: GPT2 (small, medium, large, xl), Llama3 (1B, 3B, 8B, 11B-Vision, 70B), and Gemma2 (2B, 9B, 27B) (Radford et al., 2019; Dubey et al., 2024; Team et al., 2024).
@@
Figure 1 shows our results, with most language models demonstrating high Pearson correlations. These results indicate that most language models within the same family share similar relative orientations of token embeddings. Note that this is despite the fact that the language models have different embedding dimensions. Further note that for Llama3, the base models and their instruction-tuned counterparts always have a correlation scores near 1, indicating that the global geometry of token embeddings does not change much. Lastly, in the case of untied token embeddings (Llama3 8B, 11B-V, 70B), we interestingly only observe high correlation scores for the unembedding space, but not for the embedding space.
We provide some thoughts regarding this difference in untied embeddings. Note that logits are produced by projecting the last hidden state onto the unembedding space: $\langle\mathbf{h}^{L-1},\mathcal{U}\rangle$. The fact that unembedding spaces converge across models suggests that so do the last hidden state $\mathbf{h}^{L-1}$. Put differently, token embeddings for untied models may start off in different places, but after $L$ transformer blocks, their representations across models converge. For tied models, the fact that embeddings also share similar geometry may merely be an artifact of the embeddings and unembeddings being tied. It is possible that the convergence of logits we observe is a case of neural collapse (Papyan et al., 2020; Wu & Papyan, 2024) (assuming each model family was trained on the same data), or a case of the theorized implicit geometry of next-token prediction, as studied by Zhao et al. (2024).
What drives this global similarity? We speculate that training data plays a large role, as suggested by the theory of Zhao et al. (2024). As preliminary evidence, we identify two models that share the same tokenizer, but are trained on different data. Namely, GPT-NeoX-20B (Black et al., 2022) and Olmo-7B (Groeneveld et al., 2024) are trained on The Pile (Gao et al., 2020) and Dolma (Soldaini et al., 2024), respectively. Their Pearson correlation has a score of 0.32, a significantly lower score than those within the same model family.
@@ 5.1 Locally Linear Embeddings of Tokens
Figure 2: LLE weight similarities in Llama3 unembeddings: language models share similar local structure. We compare LLE weights of token (un)embeddings across language models, where LLE weights capture the local structure of each embedding. Most token embeddings have very similar local structure across language models, as indicated by the density accumulated around high cosine similarity scores.
@@
Note that some tokens have a cosine similarity of 0. These may be undertrained tokens (Land & Bartolo, 2024) – tokens present in the tokenizer, but not in the training data (e.g., “_SolidGoldMagikarp”). Their embeddings are likely random weights assigned during initialization. Note that in high dimensions, random vectors are likely to be orthogonal. We check this hypothesis by comparing our undertrained tokens from that of Land & Bartolo (2024) – see Appendix C.
@@ 5.2 Measuring Intrinsic Dimension
We also characterize the local geometry of embeddings by defining a simple measure for intrinsic dimension (ID). To measure the ID of a token, we use $k$ of its nearest neighbors. We then run PCA on the $k$ points, and refer to the number of principal components needed to explain some threshold amount (95%) of the variance as the token’s intrinsic dimension. Varying hyperparameter values for $k$ and variance threshold yield similar trends: while the absolute ID values differ, tokens with lower ID yield semantically coherent clusters.
@@ 5.3 Low Intrinsic Dimensions Indicate Semantically Coherent Clusters
Quantitatively, we leverage ConceptNet (Speer et al., 2017) to define and measure semantic coherence. ConceptNet is a knowledge graph of 34 million nodes and edges, each node representing a concept (token) and edges representing one of 34 relation types. Namely, we design a “semantic coherence score” (SCS), and inspect the Spearman correlation between the SCS and intrinsic dimensions of 500 randomly sampled token embeddings.
To measure SCS, for each token $x$, we look at its $k$ (=50) nearest neighbors in embedding space. We then compute the average shortest distances from token $x$ to tokens $x_{k}$ in ConceptNet, with some cutoff threshold $L$ (=5):
$SCS(x)=1-\frac{1}{k}\sum_{i=0}^{k-1}\hat{d}(x,x_{i}),\quad\hat{d}(x,y)=\frac{min(d(x,y),L)}{L}$ (4)
where $d(x,y)$ is the shortest path from token $x$ to token $y$ in ConceptNet.
SCS takes a value between 0-1: 0 means that $x$ is not connected to any of its $k$ neighbors within $L$ hops, while 1 means $x$ is connected to all $k$ neighbors within $L$ hops.
Finally, for a set of tokens we measure the Spearman correlation between their SCS and intrinsic dimensions. Correlation scores and p-values are provided in Table 2. See Appendix D for scatter plots of SCS versus intrinsic dimensions.
GPT2 and Llama3 consistently demonstrates high negative correlations between SCS and intrinsic dimensions. We do not observe a correlation in Gemma2 - we conjecture that this is because of the high number of special tokens, emojis, and non-English tokens in Gemma2, which is less suitable for ConceptNet (Gemma2 has a vocabulary size of 256k, as opposed to 128k in Llama3 and 50k in GPT2).
What determines the ID of tokens? We view this as an exciting question for future work.
Model Spearman Corr. P-Value
GPT2-Medium (k=100) -0.67 1e-66
Llama3-3B (k=100) -0.45 7e-27
Gemma2-2B (k=100) 0.22 2e-7
Table 2: Spearman Correlations between Semantic Coherence Score (SCS) vs. Intrinsic Dimension High (negative) correlations indicate that tokens with low intrinsic dimensions lead to highly semantically coherent clusters. $k$ indicates the number of nearest neighbors used to calculate intrinsic dimension, not to be confused with $k$ used to measure SCS. See Table 4 for results on more hyperparameters.
@@ 6 Emb2Emb: Transferring Steering Vectors
Our setup for Llama3 is the same as that of Rimsky et al. (2023). We take human-written evaluation datasets from Perez et al. (2022) and Rimsky et al. (2023), which contain questions with two answer choices. One choice answers the question in a way that demonstrates the target behavior, while the other does not. Examples can be found in Appendix E.
@@ Results.
Figure 4 demonstrates steering Llama3-8B with steering vectors transferred from Llama3-1B and 3B (See Appendix F for more examples of transferred steering). In each subplot, the dotted curve indicates a point of reference in which we steer the target model using a steering vector from the same model (i.e., $\mathbf{v}_{S}$ == $\mathbf{v}_{T}$), while solid lines indicate a steering vector transferred from a different model. The legends indicate the source model and layer from which a steering vector is taken from, as well as the layer in the target model that is intervened on. The x-axis indicates how much each steering vector has been scaled ($\alpha$ of Equation 6), while the y-axis indicates the likelihood of the model choosing an option that exhibits a behavior of interest. In most cases, the transferred steering vectors exhibit similar trends as the original steering vector.
@@ 7 Discussion
Our findings demonstrate that the embeddings of language models share similar global and local geometric structures.
Globally, token embeddings often have similar relative orientations. Locally, token embeddings also share similar structures. We demonstrate this by comparing LLE weights of tokens, as well as by using a simple measure for intrinsic dimension. Our ID measure reveals that tokens with lower intrinsic dimensions form semantically coherent clusters.
Based on such findings, we introduce Emb2Emb, a simple application for model interpretability: steering vectors can be linearly transformed across language models.
@@ Appendix A Closed Form Solution for LLE
Appendix B Additional Results for LLE Similarity
Figures 6, 7, and 8 demonstrate LLE similarity results for GPT2, Llama3 (embeedings), and Gemma2.
Appendix C Undertrained Tokens
In Sections 5.1 and B(Figures 2, 6, 8), we observe tokens whose LLE weights have a 0 cosine similarity score across models. We hypothesize that these are undertrained tokens (Land & Bartolo, 2024), i.e., tokens that show up in the tokenizer but not in the training data. Thus these token embeddings are randomly initialized and never properly updated.
We verify our hypothesis by checking undertrained tokens identified via our method against those identified by Land & Bartolo (2024). [^2] https://github.com/cohere-ai/magikarp/tree/main
In the case of Llama3-8B, Land & Bartolo (2024) reports 2,225 undertrained tokens, while our procedure yields 13,967 candidates. Our approach matches 924 of their 2,225 tokens, yielding a recall of 0.415 and a precision of 0.066. For GPT2-Medium, Land & Bartolo (2024) reports 967 undertrained tokens, while our approach yields 3,581. Our approach matches 106 of their 967 tokens yielding a recall of 0.11 and precision of 0.029. Unfortunately, there were no undertrained tokens reported for Gemma 2.
Appendix D Spearman Correlations between Semantic Coherence Scores and Intrinsic Dimensions
Table 4 demonstrates the Spearman correlation scores between intrinsic dimensions and Semantic Coherence Scores (SCS) for a broader range of hyperparameters. Note that $k$ indicates the number of neighbors used to compute intrinsic dimension, not to be confused with $k$ used to compute SCS.
Model Spearman Corr. P-Value
GPT2-Medium (k=100) -0.67 1e-66
GPT2-Medium (k=300) -0.63 2e-56
GPT2-Medium (k=500) -0.59 6e-48
GPT2-Medium (k=1000) -0.56 5e-42
Llama3-3B (k=100) -0.45 7e-27
Llama3-3B (k=300) -0.46 1e-27
Llama3-3B (k=500) -0.37 8e-18
Llama3-3B (k=1000) -0.18 4e-5
Gemma2-2B (k=100) 0.22 2e-7
Gemma2-2B (k=300) 0.10 0.02
Gemma2-2B (k=500) 0.03 0.5
Gemma2-2B (k=1000) 0.02 0.6
Table 4: Spearman Correlations between Semantic Coherence Score (SCS) vs. Intrinsic Dimension High (negative) correlations indicate that tokens with low intrinsic dimensions lead to highly semantically coherent clusters. $k$ indicates the number of nearest neighbors used to calculate intrinsic dimension, not to be confused with $k$ used to measure SCS.
Figures 9, 10, 11 show scatter plots of intrinsic dimensions versus SCS.
Figure 9: Intrinsic dimensions versus Semantic Coherence Scores for GPT2-Medium.
Figure 10: Intrinsic dimensions versus Semantic Coherence Scores for Llama3-3B.
Figure 11: Intrinsic dimensions versus Semantic Coherence Scores for Gemma2-2B.
Appendix E Example of Steering Data
@@
Appendix F Additional Examples of Transferred Steering for Llama3